Created by Main.MatteoSolfaroli on 2007-06-29 - 16:33
Activities in sector 45 from Monday 2nd July
- CRYO
- Purging on Friday 29th evening and all the week end long;
- Refilling on Monday 2nd July;
- ElQA-TP4B and Monitoring installation
- On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB) maybe from 8am to 1pm. RESTRICTED ACCESS.
- CRYO-instrumentation
- Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
- Commissioning;
- Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
- TS/EL intervention
- Changing of the connectors of a BPM (LSS4R);
- Connection check of the instrumentation cables of BCT;
- TIM camera
- General check on Monday 2nd July;
- Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
- Safety Inspection
- AT/ACR Patrouille
- A Patrouille will be carried out on Wednesday 4th July just before the cooldown.
- AT/VAC
- Short access to LSS4R and LSS5L on Monday 2nd July.
- AB/BI
- LSS4R short intervention on the BI (capeau).
Comments:
Edit, attach file or add comment
Created by Main.MatteoSolfaroli on 2007-06-27 - 18:52
Going on activities in sector 45
- DFBAH and MB (diode)
- checked and closed by Wednesday 27th;
- Leak test on Thursday 28th;
- Purging on Friday 29th evening and all the week end long;
- Refilling on Monday 2nd July;
- On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB);
- CRYO-instrumentation
- Initialization of RF valves on Wednesday 27th or Thursday 28th;
- Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
- Commissioning;
- Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
- TS/EL intervention
- Changing of the connectors of a BPM (LSS4R);
- Placing of some Radmon cables on Wednesday 27th, Thursday 28th and Friday 29th (LSS5L);
- Connection check of the instrumentation cables of BCT;
- TIM camera
- General check on Monday 2nd July;
- Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
- Safety Inspection
- RAMSES
- Check of the equipment on Tuesday 26th and Wednesday 27th.
- CV Intervention
- Thursday 28th June from 2pm to 4pm
- QUI
- Caps removal, sleeve installation and then closure of the QUI performed on Thursday 28th.
- The new mechanism for the header B valve is being prepared and should be ready for Friday 29th or early next week.
Comments:
Edit, attach file or add comment
Created by Main.UnknownUser on 2007-06-26 - 18:47
This morning MCS has open the diode box of the dipole 3186. AT/MEL has connected an ohmmeter during the operation to follow the status of the short. At the time of opening the big flange of the diode box the resistance to ground started to change (increase). After the opening it could be possible to first see a quite big quantity of dust on the diode metallic box and second some big pieces of welding residuals.
After having disconnected the diode at the level of the half moons the short disappeared (infinite resistance) and in at the same time some debris felled down from the top part of the half moons. As you can see in the attached pictures the size of one of this residual fits perfectly between the live part and the metallic body of the bus bar duct. In addition in one picture you will see a carbon trace on the G10 insulating piece. This could be provoked by the current of 1A we injected to better localize the fault position flowing through the debris supported onto the G10 piece.
The system will be set back to the requested pressure condition (6 bars in the magnets, 2.2 bars in the DFBA's - which should be achieved by Tuesday 3rd July), when the ELQA-TP4B qualification of this faulty circuit will be repeated.

IMG_0871.jpg:

IMG_0933.jpg:

IMG_0909.jpg:

IMG_0882.jpg:
Comments:
Edit, attach file or add comment
Created by Main.UnknownUser on 2007-06-13 - 18:14
The two BSRT telescopes were Installed last Monday in LSS4.
LSS4L: The BSRTD has been connected to the BSRTM and the system successfully aligned with the laser.
LSS4R: The vacuum chamber in IP4 right side is still intact after the BGI bake-out incident and therefore connection and alignment of the BSRT right side will be done tomorrow !
J.J. Gras for the BI Team

BSRTD_C5L4_Front1.jpg:
Comments:
Edit, attach file or add comment
Created by Main.MatteoSolfaroli on 2007-06-08 - 18:42

RAT_20070608.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x06x08x18x41/RAT_20070608.pdf
Comments:
Edit, attach file or add comment
Created by Main.UnknownUser on 2007-06-04 - 00:37

RAT_20070601.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x06x04x00x35/RAT_20070601.pdf
Comments:
Edit, attach file or add comment